Calculate the real interest rate over the past 24 months using the 30 year constant maturity Treasury bond rate as the nominal interest rate and assuming that expected inflation was equal to actual inflation (based on the change in CPI). Make sure to include the Fisher Equation
Answer #1

Fisher Equation -

(1 + i) = (1 + r) (1 + π)

Here, 'π' stands for Actual Inflation Rate.

'r' stands for Real Interest Rate

'i' stands for Nominal Interest Rate.

Actual Inflation Rate for year 2019 ; 1.8%

Nominal Interest Rate for 30 year constant maturity treasury bond ; 2.89%

(1 + 0.0289) = (1+r)(1+0.018)

1.0289 = (1+r)(1.018)

1.0289/1.018 = 1+r

1.01 = 1+r

r = 1.01 - 1

r = 0.01

Hence, Real Interest Rate is 10%.